find an equation of the line containing the givin pair of points (-5,-6) (-3,-1)
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

find an equation of the line containing
the given pair of points (-5,-6) (-3,-1)
----------
slope = (-1--6)/(-3--5) = 5/2
---------------------
Form:
y = mx + b
----
Substitute for y,x,m and solve for "b":
-1 = (5/2)*-3 + b
-1 = -15/2 + b
b = 13/2
-------
Equation:
y = (5/2)x + (13/2)
=====================
Cheers,
Stan H.
==========

Answer by rfer(16322)

m=-1--6/-3--5
m=5/2
-------------
y=mx+b
-1=5/2(-3)+b
-1=-15/2+b
-2/2+15/2=b
b=13/2
b=6.5
----------------
y=5/2x+6.5